please help with the question

Solve the following initial value problem. y – 9y + 20y = 2x +€4x, y(0) = 0, y(0) = 2
Answer #1

Q. (0) -0, osy(0)=2. first let us our gtay+ 20y = 2x+64 us find ye (complementary solution) from correrponding homogenous eqfor this equation, the characteristic equation is 9²-90+20=0 912-591-49+2050 =) 91(7-5) -4 (7-5) = 0 r = 5,4 Let 9=4 & 92 = 55 4c e4x Ye + 4е [ 4хечи еч] - Y - все4*+ 1.czev. Substituting, всем +16 се4* x — 4 (А+ схеч% +de+) +20( Антвбеч») = Ч~+ Пby 5 and subtracting, Ф мм ІНpir. C + C = - 3 Бct tScy = - 4х5 оъ 4erties * * 10 с —45- 580 - 45 200 24 Lo 200 =) се 102 о |